OpenStudy (anonymous):

A square has an area of 25 square meters. What is its perimeter? a. 5 meters c. 20 meters b. 6.25 meters d. 25 meters

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

o-o

OpenStudy (beccaboo333):

Please explain direct answers

OpenStudy (anonymous):

how is it A

OpenStudy (anonymous):

is 20

OpenStudy (beccaboo333):

Explain direct answers

OpenStudy (mathstudent55):

The area of a square is the square of the length of a side. That means the side is the square root of the area. You know the area is 25 m^2. Square root of 25 m^2 will give you one side. 4 times one side is the perimeter.

OpenStudy (anonymous):

A square has 4 equal sides. So for side L, the perimeter P = 4L. But we know L^2= 25 square metres Therefore L = ±5 metres. As a negative length has no physical meaning we ignore L = -5 We insert the value of L = 5 in the equation P = 4L, and arrive at P = 20 metres

OpenStudy (anonymous):

So it's 25 @mathstudent55

OpenStudy (anonymous):

@mathstudent55 it will be 20

OpenStudy (mathstudent55):

One step at a time. A = 25 m^2 \(Side = \sqrt{25~m^2} \) What is the length of the side?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

umm 5?

OpenStudy (mathstudent55):

Correct, 5 m. Next step, since a square square has 4 sides of equal length, the perimeter is simply 4 times the length of one side. \(P = 4s = 4 \times 5 ~m = \)?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

20

OpenStudy (mathstudent55):

Correct, 20 m.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Then? i got it right

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

@livan_hedz Yes you did. Good job!

OpenStudy (anonymous):

@mathstudent55 thanks